### v3

# Large Pre-trained time series models for cross-domain Time series analysis tasks

Abstract
Large pre-trained models have been vital in recent advancements in domains like language and vision, making model training for individual downstream tasks more efficient and provide superior performance. However, tackling time-series analysis tasks usually involves designing and training a separate model from scratch leveraging training data and domain expertise specific to the task. We tackle a significant challenge for pre-training a foundational time-series model from multi-domain time-series datasets: extracting semantically useful tokenized inputs to the model across heterogenous time-series from different domains. We propose Large Pre-trained Time-series Models (LPTM) that introduces a novel method of adaptive segmentation that automatically identifies optimal dataset-specific segmentation strategy during pre-training. This enables LPTM to perform similar to or better than domain-specific state-of-art model when fine-tuned to different downstream time-series analysis tasks and under zero-shot settings. LPTM achieves superior forecasting and time-series classification results taking up to 40% less data and 50% less training time compared to state-of-art baselines. Code: www.github.com/AdityaLab/Samay
1 Introduction
Time-series analysis tasks are important well-studied problems such as forecasting (Hyndman and Athanasopoulos, 2018) and classification (Chowdhury et al., 2022) with applications in wide-ranging domains such as retail, meteorology, economics, and health. Recent works (Chen et al., 2021; Wang et al., 2022; Zeng et al., 2023) have shown the efficacy of purely data-driven deep learning models. However, most state-of-art neural models are known to be data-hungry and require substantial training data. Motivated by language and vision foundational models Bommasani et al. (2021), recent body of works build pre-trained time-series models Das et al. (2023b); Ansari et al. (2024); Jin et al. (2024); Rasul et al. (2023). These models are trained on diverse datasets from different domains during pre-training. They require less training resources and data and produce superior performance. These models can also be deployed without any training, in a zero-shot or few-shot setting.
These foundational models, however, require large amounts of data for pre-training, which is still a challenge for time-series datasets. Moreover, they do not provide consistent performance improvement across all the domains. We identify an important challenge to building a unified pre-trained model for time-series that is pre-trained on and deployed to multiple domains: representation of diverse time-series input into these models.
Most neural sequential models input time-series values for each time-step separately. However, unlike text data, each individual time stamp may not provide enough semantic meaning about local temporal patterns of the time series. To tackle this, Nie et al. (2022) proposed to segment the time series and input each segment as individual tokens to their transformer-based model. This simple segmentation method of tokenizing time-series has been used by recent pre-trained models Woo et al. (2024); Das et al. (2023b) to provide superior performance across multiple applications. However, segmenting input time-series uniformly with fixed-length segments, while simple, can be a very inflexible tokenizing method, especially when dealing with datasets from multiple domains with different set of underlying generative dynamics, sampling rate, noise, etc. For example, among two datasets, a dataset with a lower sampling rate (such as GDP time-series) may require longer segments than those with higher sampling rates to capture similar patterns in the model (such as heart sensors collecting data in milliseconds). However, note that the dynamics of the same time-series may vary with time Liu et al. (2024a). For, time intervals that are smoother with less complex dynamics, using longer segment sizes may suffice but intervals where time-series are complex and have multiple temporal patterns may require finer-grained segmentation. For example, for seasonal epidemics, time-series is smoother during the off-season and has more complex dynamics during outbreaks and times of higher incidence.
We tackle this important problem of representing diverse time-series datasets when training a pre-trained foundational model for time-series and propose Large Pre-trained Time-series Models (LPTM), a novel method for generating pre-trained models for time-series data across multiple domains. LPTM uses a simple transformer-based architecture and leverages masking-based self-supervised pre-training to train on multiple datasets from different domains. Our main contribution focuses on how we input time-series segments as tokens to the transformer. To overcome the challenges associated with segmentation on diverse datasets discussed above, we propose a novel adaptive segmentation module that segments the time-series of each domain based on how well it performs on self-supervised pre-training. The segmentation module uses a novel scoring mechanism during pre-training to identify an effective segmentation strategy for a domain. LPTM can be fine-tuned or deployed in a zero-shot setting to various forecasting and classification tasks in domains such as epidemiology, energy, economics, behavioural datasets, etc. LPTM provides performance on par with state-of-art models with lesser pre-training data, training data and fewer training steps. Our main contributions can be summarized as follows:
Multi-domain Pre-trained time-series model We propose a framework for generating foundational pre-trained models for time-series that are trained on multiple datasets across varied domains. LPTM solves the tokenization problem for cross-domain time-series data and proposes a novel adaptive segmentation module which is important to build pre-trained models for time-series similar to foundational models for text and vision.
Adaptive segmentation for cross-domain pre-training To optimally extract semantically useful information from time-series of different domains with varied dynamics and sampling rates for pre-training, we propose a novel adaptive segmentation module that learns segmentation strategy for each domain based on losses from self-supervised learning tasks.
State-of-art and efficient performance in diverse downstream time-series tasks We evaluate LPTM on downstream forecasting and classification tasks from multiple domains and observe that LPTM consistently provides performance similar to or better than previous state-of-art models usually under zero-shot evaluation as well as when fine-tuned with lesser training data and compute time. Overall, we also observe that LPTM typically requires less than 80% of training data used by state-of-art baselines to provide similar or better performance.
2 Problem Setup
Time-series analysis tasks
Our pre-trained model can be used for many time-series tasks including forecasting and classification from multiple benchmarks and domains. For a given downstream task let be the time-series dataset consisting of time series . A time-series analysis task’s goal is to predict important properties of the time-series. For example, the forecasting task involves predicting the future values whereas classification involves predicting the class label of the input time-series based on labelled training data.
Self-supervised pre-training on multi-domain datasets
The goal of our work is to learn useful knowledge and patterns from time-series datasets from time-series from different domains. Formally, we have access to time-series datasets from domains where the datasets of domain is denoted as where is the number of datasets in domain . Examples of these domains include epidemiology, energy forecasting, macroeconomics, traffic prediction, etc. The entire set of heterogenous multi-domain pre-train dataset is denoted as . In order to effectively pre-train LPTM on we formulate the problem as a set of self-supervised learning tasks on the set of pre-training datasets . During pre-training, we sample , a dataset and its domain label from and train the model on each of the self-supervised learning tasks in . The tasks in are self-supervised and do not require additional labels or other ground truth. These tasks mask patches of the input data and train the model to recover the original input.
Therefore, our problem can be formally stated as: Given a heterogeneous set of multi-domain datasets and their domain labels, we train a model leveraging SSL tasks that learns important patterns and knowledge that can be leveraged on fine-tuning the model to any time-series analysis task on any novel dataset from any of the domains .
3 Methodology
Similar to model pipelines used in NLP and vision, we first train a pre-trained model on multiple pre-training datasets . Most parameters of the pre-trained model are trained over all the datasets and tasks. However, we use a separate segmentation module for each dataset domain to capture varied sizes of segments that differ across datasets. These segments are used as tokens for a transformer model that shares the parameters across all the tasks. For each pre-trained and downstream task, we append a final linear layer on the output embeddings of the transformer to generate the final prediction. Note that during fine-tuning on downstream tasks we update the parameters of all the modules of LPTM.
3.1 Adaptive Segmentation module
Previous works input each time-step of a time-series as individual tokens or fixed-sized segments. Unlike text, individual time-steps do not typically provide any semantic meaning about the temporal patterns of the time-series. Therefore, Nie et al. (2022) proposed to segment the input time-series into uniform length segments and use each of the segments as tokens to the transformer model. Segments of fixed size are also too inflexible to capture semantics of sequences that show varied behaviour across time and across datasets. Different pre-trained datasets may have varied temporal scales, periodicity and other temporal dynamics that cannot be encompassed by a single uniform segmentation strategy. For example, epidemic time-series are usually observed weekly and may have characteristic properties like seasonality, peaks and sudden outbreaks that should be captured by segmentation. Economic time-series, in contrast, are typically captured every quarter and are more monotone with sudden anomalies and changes in data distribution. Moreover, using a uniform segmentation may not be ideal for time series with multi-scale trends with some time-stamps having denser temporal information requiring finer-graned segmentation than others. Therefore, our goal is to identify an independent segmentation strategy for each domain of time-series dataset.
Overview
LPTM automatically identifies an effective segmentation strategy for each dataset domain during pre-training. The overarching goal of the segmentation module involves identifying segments that lead to lower SSL loss. The segmentation module first scores all possible segments of the input time-series and the segments with the highest scores are then chosen as the output segments used to tokenise the time-series. The scoring function is trained such that the score of the segments will reflect how likely the chosen segments will lower the SSL loss.
Details
For a given input time-series , we pass it through a GRU to get hidden embeddings that models the temporal patterns of the input:
We then introduce a segment score function that provides a scalar score for any subsequence of the input time-series:
Intuitively, the score for a subsequence from time-stamp to denotes how good the given segment is for the dataset when optimizing for the SSL loss.
In next step, we sample subset of subsequences over the time-series that a) covers the entire input time-series, b) has a high score function value. While retrieving the optimal is an interesting combinatorial optimization problem, we generate using a simple process as follows: for each , we denote as the best segment starting from time-step . Then we generate the set of segments . To reduce the number of segments, we iteratively remove the lowest-scoring segments until we cannot remove any more segments without having time-steps not covered by any segments in the set. The final set of segments after pruning is denoted as . The segmentation procedure is summarized in Alg. 1.
To generate the token embeddings for each segment , we pass the embeddings through a self-attention layer used in transformers and aggregate the output embeddings.
Additionally, we concatenate the following features to the token embedding of each segment token to provide information regarding the position and length of the segment: (1) Positional encoding of the starting time-step of the segment defined as:
where is the dimensions of output embedding of self-attention over . (2) Positional encoding of the length of the segment . The final output of the segmentation module is a sequence where number of segments. The segments are arranged based on the ascending order of the first time-stamp of each segment. The token embeddings are fed into a stack of transformer layers similar to the encoder of the transformer Vaswani et al. (2017). The output of the transformer layers is a sequence of output embeddings denoted as .
3.2 Self-supervised learning Tasks
Pre-training on a wide range of heterogeneous datasets from multiple domains helps LPTM learn from useful patterns and latent knowledge across these domains that can be generalized to range downstream tasks on different domains. We propose two general self-supervised learning tasks motivated by pre-trained language models to enable LPTM to learn from all pre-trained datasets. We leverage a transformer model and use the segment token embeddings of the segmentation module. The two pre-training SSL tasks are Random Masking (RandMask) and Last token masking (LastMask). RandMask allows the model to extrapolate and interpolate masked segments of the input time-series. RandMask has also been explored for representation learning in previous works (Zerveas et al., 2021; Nie et al., 2022) but are applied on the same dataset as that used for training unlike our data and task-agnostic pre-training setup. Formally, we mask each input segment token with a probability of and decode the values of time-series of the masked segments from the output embeddings of the transformer. We use a simple GRU with a single hidden layer on the output embedding of the masked token to decode the values of the segment. We use mean-squared error as the loss . LastMask is similar to RandMask except we mask last fraction of the segments. This allows the model to forecast the future values of the time-series, an important task in many time-series domains.
3.3 Training details
Instance normalization
The values of the time-series of each dataset can vary widely based on the time-series domain. Therefore, as part of pre-processing, we first normalize the time-series of each dataset of pre-train datasets independently. Moreover, the data distribution and the magnitude of the time-series can vary across time. We use reversible instance normalization (REVIN) layer Kim et al. (2021). REVIN performs instance normalization on the input time-series and reverses the normalization of the output values. The normalization step is part of the neural model and gradients are calculated over the normalization and reverse normalization layers.
Training the score function
We use the loss from the SSL tasks to train the score function of the segmentation module and GRU in Eqn. 1. Since there is no direct gradient flow between the score function and the final predictions, due to the discrete nature of choosing the segments, we match the aggregated scores of all the chosen segments in to the negative logarithm of the total MSE loss of both SSL tasks:
where is the total loss of both SSL tasks. We also backpropagate over once every 10 batches to stabilize training since changing the segmentation strategy for every batch leads to unstable and inefficient training.
Linear-probing and fine-tuning
Kumar et al. (2022) showed that fine-tuning all the parameters of the pre-trained model for a specific downstream task can perform worse than just fine-tuning only the last layer (linear probing), especially for downstream tasks that are out-of-distribution to pre-trained data. To alleviate this, they suggest performing a two-stage fine-tuning process: we first perform linear probing followed by fine-tuning all the parameters.
4 Related Works
Neural models for time-series analysis
DeepAR Salinas et al. (2020) is a popular forecasting model that trains an auto-regressive recurrent network to predict the parameters of the forecast distributions. Deep Markov models Krishnan et al. (2017); Rangapuram et al. (2018); Li et al. (2021); Gu et al. (2021) model the transition and emission components with neural networks. Recent works have also shown the efficacy of transformer-based models on general time-series forecasting Oreshkin et al. (2019); Zhou et al. (2021); Chen et al. (2021); Zhou et al. (2022); Liu et al. (2021). However, these methods do not perform pre-training and are trained independently for each application domain. therefore, they do not leverage cross-domain datasets to generate generalized models that can be used for a wide range of benchmarks and tasks.
Self-supervised learning for time-series
Recent works have shown the efficacy of self-supervised representation learning for time-series for various classification and forecasting tasks in a wide range of applications such as modeling behavioral datasets Merrill and Althoff (2022); Chowdhury et al. (2022), power generation Zhang et al. (2019), health care Zhang et al. (2022). Franceschi et al. (2019) used triplet loss to discriminate segments of the same time-series from others. These works use methods such as contrastive losses Eldele et al. (2021); Yue et al. (2022) or other similarity metric-based techniques Tonekaboni et al. (2021). However, all these methods apply SSL on the same dataset that is used for training and may not adapt well to multi-domain datasets. There has been some recent works leveraging foundational models like LLMs for time-series forecasting across multiple applications. One set of works directly uses LLMs without fine-tuning to perform time-series forecasting via careful prompting Gruver et al. (2024); Jin et al. (2024); Liu et al. (2024d). Other works fine-tune LLMs specifically for time-series forecasting Zhou et al. (2023); Rasul et al. (2023); Ansari et al. (2024). The time-series representation used by these models includes using individual time-steps as input Ansari et al. (2024), converting each digit of the time-series to character embeddings to be directly used by LLMs Gruver et al. (2024); Jin et al. (2024) and uniform segmentation Das et al. (2023b); Woo et al. (2024); Zhou et al. (2023). LPTM performs superior to these methods while being 10x to 100x smaller than the large LLMs used as backbones.
5 Experiment Setup
Datasets
We derive pre-train time-series datasets from multiple domains: Epidemics: We use a large number of epidemic time-series aggregated by Project Tycho (van Panhuis et al., 2018). from 1888 to 2021 for different diseases collected at state and city levels in the US. We remove time series with missing data and use time series for 11 diseases of very diverse epidemic dynamics such as seasonality, biology, geography, etc.: Hepatitis A, measles, mumps, pertussis, polio, rubella, smallpox, diphtheria, influenza, typhoid and Cryptosporidiosis (Crypto.). Electricity: We use ETT electricity datasets (ETT1 and ETT2) collected from (Zhou et al., 2021) at 1 hour intervals over 2 years. We use the default 12/4/4 train/val/test split and use the train split for pre-training as well. Traffic Datasets: We use 2 datasets related to traffic speed prediction. PEMS-Bays (PEM-B) and METR-LA (Li et al., 2017) are datasets of traffic speed at various spots collected by the Los Angeles Metropolitan Transportation Authority and California Transportation Agencies over 4-5 months. Demand Datasets: We use bike and taxi demand datasets (NY-B, NY-T) from New York City collected from April to June 2016 sampled every 30 minutes. We all but the last 5 days of data for training and pre-training. Stock forecasting: We also collect the time-series of daily stock prices of Nasdaq and S&P 500 index using Yahoo finance data (Xu and Berkely, 2014) from July 2014 to June 2019. We train and pre-train using the first 800 trading days and use the last 400 for testing. M4 competition time-series: We also used the 3003 time-series of M4 forecasting competition (Makridakis and Hibon, 2000) which contains time-series from multiple domains including demographics, finance, and macroeconomics. Motion and behavioral sensor datasets: We use the set of sensor datasets extracted from UEA archive (Bagnall et al., 2018) and UCI Machine learning repository (Asuncion and Newman, 2007) similar to (Chowdhury et al., 2022). Note that our publicly accessible pre-training dataset is significantly smaller than other pre-trained datasets used by past work Ansari et al. (2024); Das et al. (2023b) some of which use confidential data inaccessible to us. We also do not use any synthetic datasets like Das et al. (2023b); Ansari et al. (2024).
Downstream tasks
We test the pre-trained LPTM trained on datasets discussed above on multiple forecasting and time-series classification tasks. We perform forecasting on the influenza incidence time series in US and Japan. Specifically, we use the aggregated and normalized counts of outpatients exhibiting influenza-like symptoms released weekly by CDC111https://gis.cdc.gov/grasp/fluview/fluportaldashboard.html. For influenza in Japan, we use influenza-affected patient counts collected by NIID222https://www.niid.go.jp/niid/en/idwr-e.html. We forecast up to 4 weeks ahead over the period of 2004 to 2019 flu seasons using a similar setup as Flusight competitions Reich et al. (2019).
We also perform electricity forecasting on the ETT1 and ETT2 datasets using the train/test split mentioned previously. The last 10% of PEM-Bays dataset is used for traffic forecasting up to 1 hour ahead and the last 5 days of New York demand datasets for demand forecasting up to 120 minutes in the future. We also perform forecasting on the Nasdaq dataset for up to 5 days ahead and M3 time-series for 1 month ahead. We use 6 of the sensor datasets from Asuncion and Newman (2007) for time-series classification tasks. We use an 80-20 train-test split similar to Chowdhury et al. (2022).
Baselines
We compared LPTM’s performance in various time-series tasks against twenty two state-of-the-art general forecasting and domain-specific baselines. First, we compare against recent pre-trained foundational time-series models: (1) LLM-Time Gruver et al. (2024), (2) TimesFM Das et al. (2023b), (3) Lag-LLAMA Rasul et al. (2023), (4) Chronos Ansari et al. (2024) and (5) MOIRAI Woo et al. (2024). We skip models like Time-LLM Jin et al. (2024), MOMENT Goswami et al. (2024) and Autotimes Liu et al. (2024b) which cannot perform zero-shot forecasting across domains and are outperformed by the aforementioned recent models when fine-tuned. We compared with (6) Informer Zhou et al. (2021), (7) Autoformer Chen et al. (2021), (8) iTransformer Liu et al. (2023) and (9) PatchTST Nie et al. (2022), four state-of-the-art transformer-based forecasting models. We also compare against other recent model (10) MICN (Wang et al., 2022), (11) TiDE Das et al. (2023a) (12) TFT Lim et al. (2021) and (13) TimesNeT Wu et al. (2023). We also compare with (14) N-HITS Challu et al. (2023) which uses multi-scale interpolation and (15) AutoARIMA Hyndman and Khandakar (2008) a ARIMA based model that does automatic hyperparameter search. We also compare it against three other state-of-art self-supervised methods for time-series: (16) TS2Vec (Yue et al., 2022), (17) TS-TCC (Eldele et al., 2021) and (18) SimMTM Dong et al. (2024) uses masking as pre-trained task for time-series classification.
Finally, we compared against the best models for individual tasks for each domain. For influenza forecasting, we compared against previous state-of-art models (19) EpiFNP Kamarthi et al. (2021) and (20) ColaGNN Deng et al. (2020) respectively. We also compare against (21) STEP Shao et al. (2022) a SOTA model for demand forecasting, traffic prediction, and stock prediction benchmarks among the baselines by automatically modelling sparse relations between multiple features of the time-series. For classification tasks, we compare against (22) TARNet Chowdhury et al. (2022).
6 Results
| Model | Flu-US | Flu-japan | ETT1 | ETT2 | PEM-B | NY-B | NY-T | Nasdaq | M4 |
| LLM-Time | 1.38 | 1411 | 0.57 | 0.54 | 4.3 | 4.5 | 13.53 | 0.29 | 1.189 |
| TimesFM | 1.35 | 1259 | 0.61 | 0.59 | 3.9 | 3.9 | 13.11 | 0.29 | 1.211 |
| Lag-LLAMA | 1.52 | 1488 | 0.83 | 1.06 | 5.3 | 3.8 | 12.84 | 0.24 | 1.311 |
| Chronos | 1.29 | 1274 | 0.62 | 0.56 | 4.2 | 3.6 | 13.74 | 0.29 | 1.125 |
| MOIRAI | 1.39 | 1411 | 0.69 | 0.52 | 4.2 | 4.4 | 13.82 | 0.27 | 1.192 |
| TS2Vec | 1.94 | 1233.1 | 1.33 | 1.82 | 3.7 | 4.1 | 14.39 | 0.87 | 1.616 |
| TS-TCC | 2.17 | 1356.15 | 1.14 | 1.57 | 4.1 | 3.8 | 15.72 | 0.92 | 1.492 |
| SimMTM | 2.17 | 1356.15 | 1.14 | 1.57 | 4.1 | 3.8 | 15.72 | 0.92 | 1.492 |
| LPTM | 1.14 | 996 | 0.53 | 0.49 | 3.4 | 3.2 | 13.12 | 0.22 | 0.972 |
| Model | Flu-US | Flu-Japan | ETT1 | ETT2 | PEM-B | NY-B | NY-T | Nasdaq | M4 | Rank |
| AutoARIMA | 2.14 | 1344 | 0.73 | 0.64 | 4.1 | 4.13 | 16.43 | 0.62 | 1.89 | 25.06 |
| Informer | 1.62 | 1139 | 0.57 | 0.71 | 3.1 | 2.89 | 12.33 | 0.83 | 1.055 | 15.89 |
| Autoformer | 1.41 | 1227 | 0.72 | 0.82 | 2.7 | 2.73 | 12.71 | 0.19 | 0.887 | 13.67 |
[…]
| N-HITS | 1.42 | 1211 | 0.53 | 0.62 | 2.9 | 2.74 | 11.87 | 0.57 | 0.968 | 13.0 |
| TiDE | 1.21 | 1186 | 0.49 | 0.49 | 3.5 | 3.86 | 11.95 | 0.57 | 1.078 | 13.44 |
[…]
| TimesNet | 1.04 | 1194 | 0.56 | 0.62 | 3.9 | 2.83 | 11.82 | 0.19 | 1.055 | 12.11 |
| TFT | 1.21 | 1876 | 0.52 | 0.51 | 4.6 | 2.95 | 12.55 | 0.24 | 1.18 | 16.11 |
| iTransformer | 1.14 | 1256 | 0.57 | 0.59 | 4.3 | 2.83 | 13.16 | 0.29 | 1.125 | 17.5 |
| LLM-Time | 1.21 | 1319 | 0.52 | 0.49 | 3.9 | 3.7 | 12.11 | 0.21 | 1.064 | 14.05 |
| TimesFM | 1.32 | 1214 | 0.58 | 0.49 | 3.7 | 2.8 | 12.19 | 0.22 | 1.07 | 13.44 |
| Lag-LLAMA | 1.46 | 1416 | 0.61 | 0.57 | 3.9 | 2.9 | 13.43 | 0.28 | 1.33 | 20.16 |
| Chronos | 1.21 | 1228 | 0.59 | 0.52 | 3.7 | 3.1 | 12.82 | 0.27 | 1.04 | 15.55 |
| MOIRAI | 1.31 | 1336 | 0.62 | 0.55 | 3.9 | 3.5 | 13.71 | 0.24 | 1.21 | 19.22 |
[…]
| EpiFNP | 0.52 | 872 | 0.81 | 1.25 | 4.1 | 2.98 | 12.11 | 0.28 | 1.281 | 16.77 |
| ColaGNN | 1.65 | 694 | 0.72 | 1.19 | 3.9 | 3.19 | 14.97 | 0.25 | 1.185 | 19.22 |
| TS2Vec | 1.85 | 905.9 | 0.99 | 1.74 | 3.5 | 3.11 | 13.48 | 0.94 | 1.344 | 21.94 |
| SimMTM | 1.31 | 1289 | 0.61 | 0.55 | 3.4 | 3.1 | 12.79 | 0.28 | 1.284 | 17.94 |
| TS-TCC | 1.94 | 1134.6 | 0.75 | 1.29 | 3.3 | 2.97 | 15.55 | 0.76 | 1.274 | 21 |
[…]
| LPTM-NoSegment | 0.93 | 766 | 0.57 | 0.55 | 3.2 | 3.17 | 14.96 | 0.27 | 1.146 | 13.72 |
| LPTM-NoPreTrain | 0.96 | 827 | 0.46 | 0.57 | 3.7 | 2.66 | 12.43 | 0.25 | 1.271 | 11.66 |
| LPTM-NoLinProb | 0.92 | 885 | 0.43 | 0.53 | 3.1 | 2.49 | 12.17 | 0.19 | 1.032 | 6.55 |
| LPTM-OnlyRandMask | 0.87 | 895 | 0.51 | 0.52 | 2.8 | 2.42 | 12.36 | 0.21 | 1.076 | 8.0 |
| LPTM-OnlyLastMask | 0.79 | 773 | 0.44 | 0.48 | 2.7 | 2.31 | 12.04 | 0.19 | 0.916 | 3.77 |
The code for implementation of LPTM and datasets are provided at anonymized link333www.github.com/AdityaLab/Samay and hyperparameters are discussed in the Appendix. LPTM consists of 10 layers for the transformer and overall has about 100M parameters which 2x to over 10x smaller than other pre-trained time-series models.
Zero-shot forecasting
An important benefit of foundational models in language and vision is their ability to adapt to novel tasks without any fine-tuning in a zero-shot setting Brown et al. (2020). We evaluate the zero-shot performance of LPTM and other pre-trained baselines. Similar to Gruver et al. (2024) we use the last 20% of the datasets for zero-shot evaluation. We do not fine-tune the models but only input the normalized time-series for new tasks directly during inference. The results are summarized in Table 1. LPTM outperforms all the baselines significantly. Moreover, the baselines such as TS2Vec, TS-TCC and SimMTM which are not designed to handle datasets from multiple domains perform much worse than other pre-trained methods. This shows the importance of adaptive segmentation-based tokenization of LPTM to better generalize across multiple domains.
Fine-tuned forecasting
We also train the models with training data to evaluate fine-tuned performance. The forecasting performance are shown in Table 2. We evaluated LPTM against seventeen other forecasting baselines. LPTM is either the first or a close second best-performing model in all the benchmarks. LPTM generally outperforms general time-series forecasting methods as well as 2recent pre-trained models in all benchmarks despite its much lower parameter count (2-10x lower). Further, it is competitive or superior in performance to domain-specific methods designed specifically for the given domains (such as EpiFNP and STEP). LPTM beats the previous state-of-art domain-specific baselines in five of the benchmarks and comes second in four more. Finally, LPTM improves upon the state-of-art on electricity forecasting, traffic forecasting, and M4 datasets.
Time-series classification
Unlike other autoregressive foundational models designed for forecasting, LPTM can be used for classification due to its encoder-style architecture. We add a single classification layer over pooled output embeddings to predict the class logits. We evaluate LPTM and baselines on the classification of 35 sensor and behavioral datasets from UCI classification repositiry(Asuncion and Newman, 2007). We report the accuracy scores in Table 3. We observe that LPTM has highest mean rank and largest number of times it outperforms all baselines.
| Informer | Autoformer | TimesNet | TARNet | TS2Vec | TS-TCC | TST | SimMTM | CRT | LPTM | |
| BasicMotions | 0.95 | 0.93 | 0.92 | 1.00 | 0.99 | 1.00 | 0.92 | 0.86 | 0.88 | 1.00 |
[…]
| Handwriting | 0.16 | 0.18 | 0.29 | 0.24 | 0.52 | 0.55 | 0.37 | 0.64 | 0.52 | 0.51 |
[…]
| Phoneme | 0.11 | 0.13 | 0.15 | 0.17 | 0.28 | 0.226 | 0.29 | 0.26 | 0.18 | 0.32 |
[…]
| UWaveGesture | 0.79 | 0.86 | 0.82 | 0.87 | 0.91 | 0.92 | 0.86 | 0.74 | 0.82 | 0.94 |
[…]
| SelfRegulationSCP2 | 0.562 | 0.598 | 0.612 | 0.622 | 0.593 | 0.575 | 0.604 | 0.614 | 0.625 | 0.691 |
| Occupancy | 0.774 | 0.739 | 0.814 | 0.833 | 0.876 | 0.865 | 0.881 | 0.826 | 0.814 | 0.836 |
| MosquitoSound | 0.439 | 0.493 | 0.551 | 0.632 | 0.649 | 0.662 | 0.691 | 0.703 | 0.624 | 0.715 |
Data efficiency
A significant advantage of leveraging pre-trained models is that we do not require large datasets for fine-tuning to a specific task. We evaluate the efficacy of LPTM to train with a fraction of training data. For each time-series analysis task, we fine-tune the model using only of training data for different values of . We use the first of the timestamps’ values. We do not choose a random sample to prevent data mixing from the rejected portion of training data. We also performed the similar experiment on the best baseline for each task and compare data efficiency of baseline with LPTM.
The comparison plots are shown in Figure 2. With lesser data, the performance of the baseline is much worse whereas LPTM typically requires much less data to provide similar performance to when we have access to the full dataset. This shows the importance of pre-training to quickly ramp up the performance of the model with much less data, a problem we encounter is many real-world settings such as when we need to deploy a forecasting model on novel applications such as a new pandemic with sparse data availability.
Training efficiency
An important advantage of pre-trained models is that they require much less training time and resources to fine-tune to a downstream task compared to time required for pre-training or even training from scratch. We compare the fine-tuning time of LPTM with baselines on benchmarks from different domains. We also measure the average time required by LPTM to reach the performance of the best baseline in cases where we eventually outperform them. The training times are summarized in Appendix Table 4. We observe that the time taken by LPTM to reach the performance of best best-performing baseline (LPTM-TB) is significantly smaller than the time taken by any other baselines. Even when LPTM doesn’t outperform the best baseline, it typically converges much faster.
Ablation and Sensitivity Studies
We study the impact of our adaptive segmentation as well as pre-training and linear probing via the ablation models LPTM-NoSegment, LPTM-NoPreTrain and LPTM-NoLinProb. We also investigate the individual impact of both SSL task via the ablation models LPTM-OnlyRandMask and LPTM-OnlyLastMask. The performance of the ablation variants are also shown in Tables 2. We observe that the ablation variants’ performances are significantly worse than the variants, underperforming some of the baselines. The worst performing variant is usually LPTM-NoSegment, showing the importance of deriving good time-series segments to improve representation learning of time-series for each dataset. We also examined the sensitivity of hyperparameter for SSL tasks and found the optimal value at 0.4 for LASTMASK and 0.2 for RANDMASK. The sensitivity analysis plots are in Appendix Fig. 5.
Segments generated by LPTM
We also visualized the segments in Fig. 3. We observe that the segment sizes are smaller in regions of high variance or important parts of the time-series such as peak of the epidemic whereas simpler trends have longer segments which matches our intuition.
7 Conclusion
We make a significant contribution towards general pre-trained models for time-series analysis tasks replicating the success of large pre-trained models in language and vision domains. We introduce LPTM, a general pre-trained model that provides state-of-art performance on a wide range of forecasting and classification tasks from varied domains and applications. LPTM provides similar to or better performance to state-of-art domain-specific models in applications such as epidemiology, energy, traffic, and economics and significantly beats recent time-series foundational models. We also observe that LPTM required significantly lesser pre-training and training data to reach optimal performance compared to other baselines in most benchmarks.
Our work mainly focuses on the important challenge of providing semantically meaningful inputs to the model that caters to learning time-series segmentation strategies specific to each domain. This is crucial when pre-training on diverse datasets, a key challenge for time-series data. The underlying model architecture is a straightforward transformer encoder that uses well-known masking techniques for self-supervised pre-training. Therefore, our method can be extended to leverage novel time-series model architectures and SSL methods. Extending our methods to provide calibrated forecasts that provide reliable uncertainty measures is also another important direction of research. We can also extend it to leverage multimodal datasets like text that provide important contextual information about the dataset Liu et al. (2024c).
Since our model can be applied to any generic time-series analysis tasks including those in critical domains such as public health, medicine, economics, etc., important steps need to be taken to address potential misuse of the our methods such as testing for fairness, data quality issues, ethical implications of predictions, etc.
Acknowledgements
This paper was supported in part bythe NSF (Expeditions CCF-1918770, CAREER IIS-2028586, Medium IIS-1955883, Medium IIS-2403240, Medium IIS-2106961, PIPP CCF-2200269), CDC MInD program, Meta faculty gifts, and funds/computing resources from Georgia Tech.
Appendix for Large Pre-trained time series models for cross-domain Time series analysis tasks
Appendix A Hyperparameters
The model is run on Intel Xeon CPU with 64 cores and 128 GB RAM. We use a single A100 GPU with 80GB memory. For GRU we use a single hidden layer of 50 hidden units. Dimension of is also 50. The transformer architecture consists of 10 layers with 8 attention heads each. For forecasting tasks, we train a separate decoder module with 4 more layers during fine-tuning whereas for classification we aggregate the embeddings of the last transformer layer and feed them into a single linear layer that provides logits for all classes. The SSL pre-training was done till convergence via early stopping with patience of 1000 epochs. We observed that LPTM takes 5000-8000 epochs to finish pre-training which takes around 3-4 hours. (Note that pre-training is a one-time step and downstream fine-tuning takes much less time and epochs). For both pre-training and fine-tuning, we used the Adam optimizer with a learning rate of 0.001. The hyperparameters are tuned sparingly for both LPTM and baselines from their default settings. For RandMask, we found the optimal , and for LastMask was optimal.
Appendix B Data efficiency
| Model | Flu-US | ETT2 | PEM-Bays | NY-Bike | Nasdaq | M3 | BasicMotions | EigenWorms |
| Informer | 27.3 | 25.5 | 45.1 | 49.7 | 27.1 | 49.6 | 17.5 | 14.3 |
| Autoformer | 19.5 | 29.3 | 49.5 | 55.2 | 18.5 | 45.1 | 11.9 | 19.7 |
| N-HITS | 15.4 | 22.5 | 36.1 | 49.3 | 26.4 | 47.5 | NA | NA |
| PatchTST | 12.9 | 29.5 | 36.4 | 45.7 | 18.2 | 49.4 | NA | NAg |
| MICN | 17.6 | 15.7 | 39.7 | 41.1 | 19.2 | 33.9 | NA | NA |
| TimesNet | 15.4 | 19.7 | 37.4 | 46.3 | 24.1 | 36.5 | 9.4 | 11.5 |
| STEP | 25.4 | 34.1 | 52.7 | 74.3 | 29.7 | 52.8 | NA | NA |
| EpiFNP | 22.5 | 39.5 | 41.1 | 39.1 | 21.6 | 97.6 | NA | NA |
| ColaGNN | 34.7 | 33.6 | 53.1 | 47.6 | 32.1 | 72.2 | NA | NA |
| TARNet | NA | NA | NA | NA | NA | NA | 13.7 | 9.4 |
| TS2Vec | 29.3 | 28.2 | 41.9 | 41.9 | 29.8 | 67.4 | 9.3 | 13.2 |
| TS-TCC | 21.7 | 23.7 | 46.3 | 44.3 | 25.3 | 55.8 | 12.7 | 11.1 |
| LPTM | 12.2 | 19.3 | 41.9 | 37.5 | 17.3 | 31.2 | 6.1 | 12.7 |
| LPTM-TB | NA | 12.5 | 29.6 | 32.9 | NA | 23.7 | 6.1 | 8.1 |
Appendix C Effect of SSL hyperparameter
| Model | Flu-US | Flu-Japan | ETT1 | ETT2 | PEM-Bays | NY-Bike | NY-Taxi | Nasdaq | M4 |
| AutoARIMA | 0.043180 | 23.434077 | 0.014220 | 0.035201 | 0.056066 | 0.044572 | 0.151146 | 0.025181 | 0.024060 |
| Informer | 0.044072 | 23.918021 | 0.014514 | 0.035927 | 0.057224 | 0.045492 | 0.154268 | 0.025701 | 0.024557 |
| Autoformer | 0.042310 | 22.961906 | 0.013934 | 0.034491 | 0.054936 | 0.043674 | 0.148101 | 0.024673 | 0.023575 |
| PatchTST | 0.043320 | 23.509948 | 0.014266 | 0.035314 | 0.056248 | 0.044716 | 0.151636 | 0.025262 | 0.024138 |
| N-HITS | 0.048238 | 26.178970 | 0.015886 | 0.039324 | 0.062633 | 0.049793 | 0.168850 | 0.028130 | 0.026878 |
| MICN | 0.030538 | 16.573161 | 0.010057 | 0.024895 | 0.039651 | 0.031522 | 0.106894 | 0.017808 | 0.017016 |
| TimesNet | 0.021264 | 11.539913 | 0.007003 | 0.017334 | 0.027609 | 0.021949 | 0.074431 | 0.012400 | 0.011848 |
| LLM-Time | 0.040279 | 21.859627 | 0.013265 | 0.032836 | 0.052299 | 0.041577 | 0.140991 | 0.023489 | 0.022443 |
| TimesFM | 0.034938 | 18.961321 | 0.011506 | 0.028482 | 0.045365 | 0.036065 | 0.122298 | 0.020375 | 0.019468 |
| Lag-LLAMA | 0.028790 | 15.624783 | 0.009481 | 0.023470 | 0.037382 | 0.029719 | 0.100777 | 0.016789 | 0.016042 |
| Chronos | 0.040037 | 21.728648 | 0.013185 | 0.032639 | 0.051986 | 0.041328 | 0.140146 | 0.023348 | 0.022309 |
| STEP | 0.028283 | 15.349319 | 0.009314 | 0.023056 | 0.036723 | 0.029195 | 0.099001 | 0.016493 | 0.015759 |
| EpiFNP | 0.026098 | 14.163743 | 0.008595 | 0.021275 | 0.033887 | 0.026940 | 0.091354 | 0.015219 | 0.014542 |
| ColaGNN | 0.038447 | 20.865267 | 0.012661 | 0.031342 | 0.049920 | 0.039686 | 0.134578 | 0.022421 | 0.021422 |
| TS2Vec | 0.032028 | 17.381828 | 0.010548 | 0.026109 | 0.041586 | 0.033061 | 0.112110 | 0.018677 | 0.017846 |
| SimMTM | 0.053732 | 29.160611 | 0.017695 | 0.043802 | 0.069767 | 0.055464 | 0.188081 | 0.031334 | 0.029939 |
| TS-TCC | 0.044399 | 24.095686 | 0.014622 | 0.036194 | 0.057649 | 0.045830 | 0.155414 | 0.025892 | 0.024739 |
| LPTM | 0.048677 | 26.417645 | 0.016031 | 0.039682 | 0.063204 | 0.050247 | 0.170390 | 0.028387 | 0.027123 |
| LPTM-NoSegment | 0.045970 | 24.948198 | 0.015139 | 0.037475 | 0.059689 | 0.047452 | 0.160912 | 0.026808 | 0.025614 |
| LPTM-NoPreTrain | 0.058905 | 31.968143 | 0.019399 | 0.048020 | 0.076484 | 0.060804 | 0.206190 | 0.034351 | 0.032822 |
| LPTM-NoLinProb | 0.025968 | 14.092985 | 0.008552 | 0.021169 | 0.033718 | 0.026805 | 0.090898 | 0.015143 | 0.014469 |
| LPTM-OnlyRandMask | 0.039517 | 21.446157 | 0.013014 | 0.032214 | 0.051310 | 0.040791 | 0.138324 | 0.023045 | 0.022019 |
| LPTM-OnlyLastMask | 0.058067 | 31.513665 | 0.019123 | 0.047337 | 0.075397 | 0.059940 | 0.203258 | 0.033863 | 0.032355 |
| Model | Score |
| AutoARIMA | 21.388889 |
| Informer | 14.055556 |
| Autoformer | 12.000000 |
| PatchTST | 6.833333 |
| N-HITS | 11.388889 |
| MICN | 7.277778 |
| TimesNet | 11.111111 |
| LLM-Time | 12.777778 |
| TimesFM | 12.277778 |
| Lag-LLAMA | 17.444444 |
| Chronos | 13.888889 |
| STEP | 9.111111 |
| EpiFNP | 14.666667 |
| ColaGNN | 16.500000 |
| TS2Vec | 19.111111 |
| SimMTM | 15.611111 |
| TS-TCC | 18.111111 |
| LPTM | 2.500000 |
| LPTM-NoSegment | 12.055556 |
| LPTM-NoPreTrain | 10.444444 |
| LPTM-NoLinProb | 6.222222 |
| LPTM-OnlyRandMask | 7.555556 |
| LPTM-OnlyLastMask | 3.666667 |
NeurIPS Paper Checklist
1.
Claims
Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
Answer: [Yes]
Justification: The contributions listed in the Introduction are clearly expanded in Methodology and Results
Guidelines:
The answer NA means that the abstract and introduction do not include the claims made in the paper.
The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
2.
Limitations
Question: Does the paper discuss the limitations of the work performed by the authors?
Answer: [Yes]
Justification: Conclusions discuss limitiations such as not applicable to probabilistic and multivariate forecasts and potential extensions as future work.
Guidelines:
The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
The authors are encouraged to create a separate "Limitations" section in their paper.
The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
3.
Theory Assumptions and Proofs
Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
Answer: [N/A]
Justification: No theoretical results.
Guidelines:
The answer NA means that the paper does not include theoretical results.
All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
All assumptions should be clearly stated or referenced in the statement of any theorems.
The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
Theorems and Lemmas that the proof relies upon should be properly referenced.
4.
Experimental Result Reproducibility
Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
Answer: [Yes]
Justification: We provide model architecture and training details in methodology, link to the anonymized repo and hyperparameters in Appendix A.
Guidelines:
The answer NA means that the paper does not include experiments.
If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
5.
Open access to data and code
Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
Answer: [Yes]
Justification: Data and code available in https://github.com/AdityaLab/LPTM/
Guidelines:
The answer NA means that paper does not include experiments requiring code.
Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
6.
Experimental Setting/Details
Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
Answer: [Yes]
Justification: Specific details on hyperparameters are in Appendix A. Training details are in Section 3.3. Experiment setup, dataset splits are in Sec 5.
Guidelines:
The answer NA means that the paper does not include experiments.
The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
The full details can be provided either with the code, in appendix, or as supplemental material.
7.
Experiment Statistical Significance
Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
Answer: [Yes]
Justification: We perform multiple (10) runs of each experiment and report the average scores. We highlight the instances where LPTM in statistically significantly better than all baselines.
Guidelines:
The answer NA means that the paper does not include experiments.
The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
The assumptions made should be given (e.g., Normally distributed errors).
It should be clear whether the error bar is the standard deviation or the standard error of the mean.
It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
8.
Experiments Compute Resources
Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
Answer: [Yes]
Justification: We mention the compute resources used in Appendix A and training time in Appendix B.
Guidelines:
The answer NA means that the paper does not include experiments.
The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
9.
Code Of Ethics
Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
Answer: [Yes]
Justification: We went through and confirm the paper follows all the statements in the code of ethics.
Guidelines:
The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
10.
Broader Impacts
Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
Answer: [No]
Justification: Our work is a general purpose time-series model that can be utilized towards wide range of applications. Care should be taken to assess the ethics of the decision taken by the predictions, clean the data for any sensitive information and ensure fairness and positive utility in the resulting outcome.
Guidelines:
The answer NA means that there is no societal impact of the work performed.
If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
11.
Safeguards
Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
Answer: [No]
Justification: Since this work is a general-purpose foundational work on time-series and we do not focus on any potentially sensitive applications, providing safeguards for release of our work is not applicable.
Guidelines:
The answer NA means that the paper poses no such risks.
Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
12.
Licenses for existing assets
Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
Answer: [Yes]
Justification: All assets in code are either created by authors or are derived from open-source works and are cited appropriately.
Guidelines:
The answer NA means that the paper does not use existing assets.
The authors should cite the original paper that produced the code package or dataset.
The authors should state which version of the asset is used and, if possible, include a URL.
The name of the license (e.g., CC-BY 4.0) should be included for each asset.
For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
13.
New Assets
Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
Answer: [Yes]
Justification: The code is the only original asset which is released as an anonymous repo linked in the paper. We will make a public release of this on acceptance.
Guidelines:
The answer NA means that the paper does not release new assets.
Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
The paper should discuss whether and how consent was obtained from people whose asset is used.
At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
14.
Crowdsourcing and Research with Human Subjects
Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
Answer: [N/A]
Justification: No crowdsourcing research.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
15.
Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
Answer: [N/A]
Justification: This work doesn’t deal with research requiring IRB review.
Guidelines:
The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.